a chemist has one solution that is 20% alcohol and another that is 60% alcohol. How much of each should the chemist use to get 100 ml of a solution that is 52% alcohol?
Answer by stanbon(75887)

a chemist has one solution that is 20% alcohol and another that is 60% alcohol. How much of each should the chemist use to get 100 ml of a solution that is 52% alcohol?
--------------
Equation:
alcohol + alcohol = alcohol
0.20x + 0.60(100-x) = 0.52*100
---------
20x + 60*100 - 60x = 52*100
-40x = -8*100
x = 20 ml (amt of 20% solution needed)
-----
100-x = 80 ml (amt of 60% solution needed)
==============================================
Cheers,
Stan H.
================